# A short character sum in $\mathbb{F}_{p^{3}}$

### v1

@@ 4. Multiplicative energy of intervals in $\mathbb{F}_{p^{d}}$
$A_{s}$ $\leq|Q_{0}|^{2}E(I_{s},\ldots,k_{s}-times,z+Q)$
$\leq|Q_{0}|^{2}exp\left(c_{k_{s}}\frac{\log p}{\log\log p}\right).E(\mathbb{F}_{p},z+Q).$
If $z=a_{0}+a_{1}w+a_{2}w^{2}+a_{3}w^{3}+\ldots+a_{d-1}w^{d-1}$ with $a_{i}\in\mathbb{F}_{p}$ and let $Q=K\omega^{n_{0}}+K\omega^{n_{1}}+\ldots+K\omega^{n_{d-2}}$ with $K=[1,p^{\rho}]$. Then
$E(\mathbb{F}_{p},Q+z)$
$=$ |$|\{(t_{1},t_{2},\zeta_{1},\zeta_{2})\in\mathbb{F}_{p}^{2}\times Q^{2}:t_{1}(\zeta_{1}+z)=t_{2}(\zeta_{2}+z)\neq 0\}|$
$=$ |$|\{(t_{1},t_{2},x_{0},\ldots,x_{d-2},x_{0}^{\prime},\ldots,x_{d-2}^{\prime})\in\mathbb{F}_{p}^{2}\times K^{2(d-1)}:t_{1}a_{i}=t_{2}a_{i}\forall i\neq n_{j}\text{ and}$
$t_{1}((x_{0}+a_{n_{0}})\omega^{n_{0}}+\ldots+(x_{d-2}+a_{n_{d-2}})\omega^{n_{d-2}})=t_{2}((x^{\prime}_{0}+a_{n_{0}})\omega^{n_{0}}+\dots+(x^{\prime}_{d-2}+a_{n_{d-2}})\omega^{n_{d-2}})(\neq 0)\}|$
$\leq$ |$|\{(t_{1},t_{2},x_{0},\ldots,x_{d-2},x_{0}^{\prime},\ldots,x_{d-2}^{\prime})\in\mathbb{F}_{p}^{2}\times K^{2(d-1)}$
$:t_{1}(x_{i}+a_{n_{i}})=t_{2}(x_{i}+a^{\prime}_{n_{i}})(\neq 0\text{ for at least one $i$)}\}|$
@@
(3) $\begin{cases}t_{1}(x_{0}+a_{n_{0}})=t_{2}(x_{0}+a^{\prime}_{n_{0}})\\ t_{1}(x_{1}+a_{n_{1}})=t_{2}(x_{1}+a^{\prime}_{n_{1}})\\ \ldots\\ t_{1}(x_{d-2}+a_{n_{d-2}})=t_{2}(x_{d-2}+a^{\prime}_{n_{d-2}})\\ \end{cases}$
With at least one of the $t_{1}(x_{i}+a_{n_{i}})=t_{2}(x_{i}+a^{\prime}_{n_{i}})\neq 0.$ Let us assume that $t_{1}(x_{0}+a_{n_{0}})=t_{2}(x_{0}+a^{\prime}_{n_{0}})\neq 0$. Then the above equations(3) becomes
(4) $\begin{cases}t_{1}=t_{2}\\ \frac{x_{i}+a_{n_{i}}}{x_{0}+a_{n_{0}}}=\frac{x_{i}^{\prime}+a_{n_{i}}}{x_{0}^{\prime}+a_{n_{0}}},\text{ for }i=1,\ldots,d-2\\ \end{cases}$
For i=1, the number of solutions $(x_{0},x_{0}^{\prime},x_{1},x_{1}^{\prime})$ in (4) is bounded by $E([1,p^{\sigma}]+a_{0},[1,p^{\sigma}]+a_{1})$, which is bounded by $p^{2\sigma}\log p$, by [9] . Also the choice of $t\in\mathbb{F}_{p}$ is bounded by $p$. And each of the $x_{2},x_{3},\ldots,x_{d-2}$ is bounded by $p^{\sigma}$. By our assumption $t_{1}(x_{0}+a_{n_{0}})\neq 0$. So, it also requires one to count for the cases where $t_{1}(x_{1}+a_{n_{1}}),t_{1}(x_{2}+a_{n_{2}})\neq 0$ and so on in a similar way. Therefore,
(5) $E(\mathbb{F}_{p},z+Q)\leq dp^{1+\sigma(d-1)}\log p.$
and
(6) $A_{s}\leq|Q_{1}|^{2}exp(c_{k_{s}}\frac{\log p}{\log\log p}).p^{1+\sigma(d-1)}$
Let us remind that $|Q|=p^{(d-1)\sigma}$ and $|Q_{0}|=p^{\frac{1}{10}(9-4\sigma)}.$ Combining the equations 1,2 and 6 we get the following:
$E(I_{1},I_{2},\ldots,I_{r},z+Q,Q_{0})$
@@ 5. Proof of Main Theorems
5.1. Proof of Theorem 1
Theorem 1.
Given $\varepsilon>0$, there is $\delta(\varepsilon)>0$ such that if $\omega\in\mathbb{F}_{p^{3}}-\mathbb{F}_{p}$ and $I,J$ are intervals of size $p^{3/8+\varepsilon}$($p$ sufficiently large), then
$\left|\sum_{x\in I\\ y\in J}\chi(x+\omega y)\right|<p^{-\delta(\varepsilon)}|I||J|$
for $\chi$ a nontrivial multiplicative character in $\mathbb{F}_{p^{3}}$.
Proof.
Let us denote $I_{0}=[1,p^{3/8}]$ and $K=[1,p^{\lambda}],$ where $\lambda>0$ and
$\rho>\frac{3}{8}+\lambda.$
We translate $I+\omega J$ by $KK(I_{0}+\omega I_{0})$ and estimate
$\frac{1}{|K|^{2}|I_{0}|^{2}}\sum_{\begin{subarray}{c}x_{1},y_{1}\in I_{0}\\ s\in K\\ x\in I,y\in J\end{subarray}}\left|\sum_{t\in K}\chi\left(x+\omega y+st(x_{1}+\omega y_{1})\right)\right|$
(7) $=\frac{1}{|K|^{2}|I_{0}|^{2}}\sum_{\begin{subarray}{c}x_{1},y_{1}\in I_{0}\\ s\in K\\ x\in I,y\in J\end{subarray}}\left|\sum_{t\in K}\chi\left(t+\frac{x+\omega y}{s(x_{1}+\omega y_{1})}\right)\right|.$
Let $\omega(\mu)=|\{(x,y,u,v)\in I\times J\times M:\mu=\frac{x+\omega y}{u+\omega v}\}$|.
Then the above double sum becomes after using Hölder’s inequality:
$\sum_{\mu\in\mathbb{F}_{p^{3}}}\omega(\mu)\left|\sum_{t\in T}\chi(t+\mu)\right|\leq\left(\sum_{\mu\in\mathbb{F}_{p^{3}}}\omega(\mu)^{\frac{2k}{2k-1}}\right)^{1-\frac{1}{2k}}\left(\sum_{\mu\in\mathbb{F}_{p^{3}}}\left|\sum_{t\in T}\chi(\mu+t)\right|^{2k}\right)^{\frac{1}{2k}}.$
Now using Hölder’s ineqality one more time we get
(8) $\sum_{\mu\in\mathbb{F}_{p^{3}}}\omega(\mu)\left|\sum_{t\in T}\chi(t+\mu)\right|\leq\underbrace{\left[\sum_{\mu\in\mathbb{F}{p^{3}}}\omega(\mu)\right]^{1-\frac{1}{k}}}_{\alpha}\underbrace{\left[\sum_{\mu\in\mathbb{F}_{p^{3}}}\omega(\mu)^{2}\right]^{\frac{1}{2k}}}_{\beta}\underbrace{\left(\sum_{\mu\in\mathbb{F}_{p^{3}}}\left|\sum_{t\in I}\chi(\mu+t)\right|^{2k}\right)^{\frac{1}{2k}}}_{\gamma}.$
$\alpha$ is trivially bounded by
(9) $\alpha\leq(|I_{0}|^{2}|K||I||J|)^{1-\frac{1}{k}}.\$
By the definition of multiplicative energy the estimation of $\beta$ is
(10) $\beta\leq E(K,I_{0}+\omega I_{0},I+\omega J)^{1/2k}.$
By the Weil’s bound one can check a bound of $\gamma$ (See Appendix)
(11) $\gamma\leq|K|^{1/2}p^{\frac{3}{2k}}+|K|p^{\frac{3}{4k}}.$
Claim:
$E(K,I_{0}+\omega I_{0},I+\omega J)<p^{-\tau}|I|^{2}|J|^{2}|K|^{2}.$
Assuming the claim, fixing $\lambda=3/2k$ and combining equations 9,10 and 11 in 8, we get the 7 is bounded by $|I||J|p^{-\tau/2k}$. So the theorem is proved if we choose $\delta<\tau/2k.$
Proof of the claim:
$E(K,I_{0}+\omega I_{0},I+\omega J)$
$=$ |$\left|\{(k_{1},i_{0}^{\prime}+\omega i_{0}^{\prime\prime},i_{1}+\omega j_{1},k_{2},i_{0}^{\prime\prime\prime}+\omega i_{0}^{\prime\prime\prime\prime},i_{2}+\omega j_{2}):k_{1}(i_{0}^{\prime}+\omega i_{0}^{\prime\prime})(i_{1}+\omega j_{1})=k_{2}(i_{0}^{\prime\prime\prime}+\omega i_{0}^{\prime\prime\prime\prime})(i_{2}+\omega j_{2})\}\right|$
$\leq$ |$|\{(k_{1},i_{0}^{\prime},i_{0}^{\prime\prime},i_{1},j_{1},k_{2},i_{0}^{\prime\prime\prime},i_{0}^{\prime\prime\prime\prime},i_{2},j_{2}):k_{1}i_{0}^{\prime}i_{1}=k_{2}i_{0}^{\prime\prime\prime}i_{2},k_{1}i_{0}^{\prime\prime}j_{1}=k_{2}i_{0}^{\prime\prime\prime\prime}j_{2}\}|$
$\leq$ $E(K^{\prime},S^{\prime},T^{\prime})$
Where,
$S^{\prime}=\{(k,k):k\in K\}$
$T^{\prime}=\{(i_{0},i_{0}^{\prime}):i_{0},i_{0}^{\prime}\in I\}$
$R^{\prime}=\{(i,j):i\in I,j\in J\}$
Let,
$R=\{(x-\lambda_{1}y,x-\lambda_{2}y):x\in I,y\in J\}$
$T=\{(x_{0}-\lambda_{1}y_{0},x_{0}-\lambda_{2}y_{0}):x_{0},y_{0}\in I_{0}\}$
$S=\{(s,s):s\in K\}.$
Now, we define three invertible maps: $\phi_{1}:T^{\prime}\mapsto T,\phi_{2}:R^{\prime}\mapsto R,id:S^{\prime}\mapsto S$ by the following:
$\phi_{1}(x_{1},x_{2})=(x_{1}-\lambda_{1}x_{2},x_{1}-\lambda_{2}x_{2}),$
$\phi_{2}(x,y)=(x-\lambda_{1}y,x-\lambda_{2}y),$
and
$id(k,k)=(k,k).$
Note that above three maps are bijective linear correspondences. So, the corresponding multiplicative energy should be the same
$E(R^{\prime},T^{\prime},S^{\prime})=E(R,S,T).$
Now, we shall cut the things short by the estimate of [(6.7),[6]] which says
$E(R,T,S)\leq p^{-\tau}|I|^{2}|J|^{2}|K|^{2}.$
Hence, the claim is established. ∎
5.2. Proof of Theorem 2
Theorem 2.
Let $\chi$ be a nontrivial character of $\mathbb{F}_{p^{d}}$and $\omega\in\mathbb{F}_{p^{d}}$ be a generator over $\mathbb{F}^{*}_{p}$. Given $\varepsilon>0$, there exists a $\delta(\varepsilon)$ such that if $J_{0},J_{1},\ldots,J_{d-2}$ be intervals of size $p^{\rho^{\prime\prime}_{d}+\varepsilon}$, where
Then
$\sum_{q\in Q}\chi(q)<p^{-\delta(\varepsilon)}|J_{0}||J_{1}|\ldots|J_{d-2}|$
where, $Q=\{x_{0}\omega^{n_{0}}+x_{1}\omega^{n_{1}}+x_{2}\omega^{n_{2}}+\ldots+x_{d-2}\omega^{n_{d-2}}:x_{i}\in J_{i}\text{ }\forall i\in\{0,1,\ldots,d-2\},n_{0}=0<n_{1}=1<\ldots<n_{d-2}<d\}.$ and $\delta=\delta(\varepsilon)$ is uniform over $\omega$.
@@ 5.2. Proof of Theorem 2
Let us choose $k_{1},\ldots,k_{r}\in\mathbb{N}$, satisfy
@@
and carry out the Burgess argument. Now it is enough to estimate
@@
(13) $\beta\leq$ $E(Q,Q_{0},I_{1},I_{2},\ldots,I_{r})^{\frac{1}{2k}}$
@@
[…]
@@ 6. Application to Character Sums of Binary cubic form
Theorem 3.
Given $\varepsilon>0$, there is $\delta>0$ such that the following holds. Let $p$ be a large prime and $f(x,y)=x^{3}+ax^{2}y+bxy^{2}+cy^{3}$ which is not a perfect cube in moduli p. Let $I,J\subseteq[1,p-1]$ be interval of size
$|I|,|J|>p^{\frac{3}{8}+\varepsilon}$
@@
for a non principle multiplicative character $\chi(\bmod{p)}$. This estimate is uniform in $f$.
@@
Here there are four cases.
Case 1:
$f$ is irreducible $(\bmod{p})$. Then $\chi(f(x,y))$ is a character $\bmod{p}$ of $x+\omega y$ with $\omega$ is a root of the polynomial $g(x):=x^{3}+ax^{2}+bx^{2}+cx+d$ in the cubic extension $Q(\omega)$ and the result then follows from Theorem 1.
Case 2:
$f(x,y)$ is reducible in $\mathbb{F}_{p}[x,y]$ where
@@
$I_{0}=[1,\frac{1}{10}p^{3/8}]$ and $K=[1,p^{\kappa}]$. We translate $(x,y)$ by $(stx_{0},sty_{0})$ with $x_{0},y_{0}\in I_{0}$ and $s,t\in K$ and estimate
@@
(17) $w(z_{1},z_{2}):=\\ \left|\left\{(x,y,x_{0},y_{0},s)\in I\times J\times I_{0}\times I_{0}\times K:z_{1}=\frac{x-\lambda_{1}y}{s(x_{0}-\lambda_{1}y_{0})},z_{2}=\frac{x-\omega_{2}y}{s(x_{0}-\omega_{2}y_{0})}\right\}\right|.\$
@@
(18) $\frac{1}{|K|^{2}|I_{0}|^{2}}\sum\limits_{\begin{subarray}{c}z_{1}\in\mathbb{F}_{p}\\ z_{2}\in\mathbb{F}_{p^{2}}\end{subarray}}w(z_{1},z_{2})\left|\sum_{t\in K}\chi((t+z_{1})(t+z_{2})(t+z_{2}^{\prime}))\right|\$
where $z_{2},z_{2}^{\prime}$ share the same monic polynomial over $\mathbb{F}_{p}$. We shall use Holder’s inequality subsequently Weil’s Theorem to estimate the above. Using Holder’s inequality twice it is obtained
@@
The third term is estimated by using Weil’s bound[12] which we will see in appendix as Lemma 13
$\left(\sum\limits_{\begin{subarray}{c}z_{1}\in\mathbb{F}_{p}\\ z_{2}\in\mathbb{F}_{p^{2}}\end{subarray}}\left|\sum_{t\in K}\chi((t+z_{1})(t+z_{2})(t+z_{2}^{\prime}))\right|^{2k}\right)^{\frac{1}{2k}}$
$\lesssim\left(\sum\limits_{\begin{subarray}{c}z_{1}\in\mathbb{F}_{p}\\ a\in\mathbb{F}_{p}\\ b\in\mathbb{F}_{p}\end{subarray}}\left|\sum_{t\in K}\chi((t+z_{1})(t^{2}+at+b))\right|^{2k}\right)^{\frac{1}{2k}}$
(21) $\lesssim r|K|^{\frac{1}{2}}p^{\frac{3}{2r}}+|K|p^{\frac{3}{4r}}\$
@@
(22) $\sum w(z_{1},z_{2})^{2}<|I|^{2}|J|^{2}|K|^{2}p^{-\tau}$
for some $\tau>0$. Choosing $\kappa=3/2k$, we obtain an upper bound for the equation (16) of the form $|I||J|p^{-\tau/2k}$ from the above three estimates. For estimating second term we may assume $|I|,|J|<p$ here. Let
@@
(23) $E(R,T,S)\leq|I|^{2}|J|^{2}|K|^{2}p^{-\tau}$
and for that we need to calculate the multiplicative energy of the triple product of the set $(R,T,S)$. First we estimate the following:
@@
$E(R,T)<exp\left(c\frac{\log p}{\log\log p}\right).$
We shall prove the lemma in the due course. Let us start with writing $R$ as union of translates of a $T$ as bellow:
@@
$E(R,T)\leq\frac{|R|^{2}}{|T|^{2}}\max_{\zeta\in\mathbb{F}_{p}\times\mathbb{F}_{p^{2}}\times\mathbb{F}_{p^{2}}}E(T+\zeta,T).$
Thus, as in [6], it is enough to show
$\max_{\zeta,\Theta\in\mathbb{F}_{p}\times\mathbb{F}_{p^{2}}\times\mathbb{F}_{p^{2}}}E(T+\zeta,T+\Theta)<exp\left(c\frac{\log p}{\log\log p}\right)|T|^{2}.$
We can prove a more general lemma here.
@@
$\max_{\rho\in\mathbb{F}_{p}^{*}\times\mathbb{F}_{p^{2}}^{*}\times\mathbb{F}_{p^{2}}^{*}}\left|\left\{(z_{1},z_{2})\in T\times T:\rho=z_{1}z_{2}\right\}\right|<exp\left(c\frac{\log p}{\log\log p}\right)$
@@
Let us fix an $\rho=(\rho_{1},\rho_{2},\rho_{3})\in\mathbb{F}_{p}^{*}\times\mathbb{F}_{p^{2}}^{*}\times\mathbb{F}_{p^{2}}^{*}$ and consider a product representation of $\rho=z_{1}\times z_{2}$ where $z_{1},z_{2}\in T$. Then $z_{1}=(x_{1}-\lambda_{1}y_{1},x_{1}-\omega_{2}y_{1},x_{1}-\omega_{3}y_{1})$ and $z_{2}=(x_{2}-\lambda_{1}y_{2},x_{2}-\omega_{2}y_{2},x_{2}-\omega_{3}y_{2})$ with $x_{1},x_{2},y_{1},y_{2}\in I_{0}$, we want to obtain an estimate of the number of solutions of the following:
(24) $\begin{cases}(x_{1}-\lambda_{1}y_{1})(x_{2}-\lambda_{2}y_{2})=\rho_{1}(\text{ in }F_{p}(\omega))\\ \\ (x_{1}-\omega_{2}y_{1})(x_{2}+\omega_{2}y_{2})=\rho_{2}(\text{ in }F_{p}(\omega))\\ \\ (x_{1}-\omega_{3}y_{1})(x_{2}+\omega_{3}y_{2})=\rho_{2}(\text{ in }F_{p}(\omega))\end{cases}$
where $\omega$ is a root of the polynomial $x^{2}+\lambda_{2}x+\lambda_{3}$ which is $\omega_{2}$ and $\omega_{3}$. Let $\mathcal{S}$ be the solution set of the above equation. Let $(x_{1},x_{2},y_{1},y_{2}),(x_{1}^{\prime},x_{2}^{\prime},y_{1}^{\prime},y_{2}^{\prime})\in\mathcal{S}$. So, it satisfies the equation (24). So, all three of $\lambda_{1},\omega_{1},\omega_{3}$ satisfies the quadratic polynomial equation
@@
(25) $\begin{cases}(y_{1}y_{2}-y_{1}^{\prime}y_{2}^{\prime})=0\text{ in }F_{p^{2}},\\ (x_{1}x_{2}-x_{1}^{\prime}x_{2}^{\prime})=0\text{ in }F_{p^{2}}.\\ \end{cases}$
Also in the above congruences(25) the elements are in $\mathbb{F}_{p}$. So. this gives the same congruences in the field $\mathbb{F}_{p}$. Now by the definition of $I_{0}$ it’s elements are bounded by $\frac{p^{3/8}}{10}$ so $(x_{1}x_{2}-x_{1}^{\prime}x_{2}^{\prime}),(y_{1}y_{2}-y_{1}^{\prime}y_{2}^{\prime})$ are bounded by $\frac{1}{25}p^{3/4}$. Also they are elements of $\mathbb{Z}$. And Hence in $\mathbb{Z}$
@@
$y_{1}y_{2}=y_{1}^{\prime}y_{2}^{\prime}$
Finally we shall use divisor estimate $exp\left(c\frac{\log p}{\log\log p}\right)$ to get the required result. ∎
Now we shall continue with proving the equation (23). Let $\kappa=3/2k$ in the definition of $K$. Thus
@@
(26) $E(R,T,S)\leq Q^{\frac{1}{k}}.exp\left(c\frac{\log p}{\log\log p}\right).|R|^{2(1-\frac{1}{k})}$

### v3

@@ 3. Multiplicative energy of intervals in $\mathbb{F}_{p^{d}}$
[…]
Let $Q=K+K\omega+\ldots+K\omega^{d-2}$ with $K=[1,p^{\rho}]$. Then
$E(\mathbb{F}_{p},Q)$
$=$ $|\{(t_{1},t_{2},\zeta_{1},\zeta_{2})\in\mathbb{F}_{p}^{2}\times Q^{2}:t_{1}\zeta_{1}=t_{2}\zeta_{2}\neq 0\}|$
$=$ $|\{(t_{1},t_{2},x_{0},\ldots,x_{d-2},x_{0}^{\prime},\ldots,x_{d-2}^{\prime})\in\mathbb{F}_{p}^{2}\times K^{2(d-1)}:t_{1}a_{i}=t_{2}a_{i}\forall i\neq n_{j}\text{ and}$
$t_{1}(x_{0}\omega^{n_{0}}+\ldots+x_{d-2}\omega^{n_{d-2}})=t_{2}(x^{\prime}_{0}\omega^{n_{0}}+\dots+x^{\prime}_{d-2}\omega^{n_{d-2}})$
$(\neq 0)\}|$
$\leq$ $|\{(t_{1},t_{2},x_{0},\ldots,x_{d-2},x_{0}^{\prime},\ldots,x_{d-2}^{\prime})\in\mathbb{F}_{p}^{2}\times K^{2(d-1)}$
$:t_{1}x_{i}=t_{2}x_{i}^{\prime}\neq 0\text{ (for at least one $i$)}\}|$
@@
[…]
With at least one of the $t_{1}x_{i}=t_{2}x_{i}^{\prime}\neq 0.$ Let us assume that $t_{1}x_{0}=t_{2}x_{0}^{\prime}\neq 0$. Then the above equations (6) becomes
[…]
For i=1, the number of solutions $(x_{0},x_{0}^{\prime},x_{1},x_{1}^{\prime})$ in (7) is bounded by $E([1,p^{\sigma}],[1,p^{\sigma}])$, which is bounded by $p^{2\sigma}\log p$, by [11]. Also the choice of $t\in\mathbb{F}_{p}$ is bounded by $p$ and each of the $x_{2},x_{3},\ldots,x_{d-2}$ is bounded by $p^{\sigma}$. Therefore,
(8) $E(\mathbb{F}_{p},Q)\leq dp^{1+\sigma(d-1)}\log p.$
and using the above estimation in the equation 5 we obtain
(9) $A_{s}\leq|Q_{1}|^{2}\exp(c_{k_{s}}\frac{\log p}{\log\log p}).p^{1+\sigma(d-1)}.$
Let us remind that $|Q|=p^{(d-1)\sigma}$ and $|Q_{0}|=p^{\frac{1}{10}(9-4\sigma)}.$ Combining the equations 3, 4 and 9 we get the following:
$E(I_{1},I_{2},\ldots,I_{r},Q,Q_{0})$
@@ 4. Proof of main theorems
We shall directly prove the general theorem which is Theorem 2.
Proof of Theorem 2:
@@
Let us choose $k_{1},\ldots,k_{r}\in\mathbb{N}$ which satisfy
@@
and proceed with the Burgess argument as above. Now it is enough to estimate
@@
(11) $\beta=$ $E(Q,Q_{0},I_{1},I_{2},\ldots,I_{r})^{\frac{1}{2k}}$
@@
Note: $\rho_{3}=3/8=.375,\rho_{4}=.417,\rho_{5}=.438,\rho_{6}=.447\text{ and so on.}$ Also, we can observe that $\rho_{d}<1/2$. So, it is indeed a short-character sum over intervals.
Proof of Theorem 1 comes as a corollary of Theorem 2 if we take $d=3.$
@@ 5. Application to character sums of binary cubic form
Theorem.
Given $\varepsilon>0$, there is $\delta>0$ such that the following holds. Let $p$ be a large prime and $f(x,y)=x^{3}+ax^{2}y+bxy^{2}+cy^{3}$ is non-degenerate over $\mathbb{F}_{p}$. Let $I,J\subseteq[1,p-1]$ be interval of size
$p^{\frac{1}{2}}>|I|,|J|>p^{\frac{3}{8}+\varepsilon}$
@@
for any non principle multiplicative character $\chi(\bmod{p)}$. This estimate is uniform in $f$.
@@
We now distinguish three non-degenarate cases for $f(x,y)$. Here by non-degenerate cubic form we mean that the de-homogenized polynomial of $f(x,y)$ cannot have a root of multiplicity greater than $1$.
Case 1. $f$ is irreducible $(\bmod{p})$. Then $\chi(f(x,y))$ is a character $\bmod{p}$ of $x+\omega y$ with $\omega$ is a root of the polynomial $g(x):=x^{3}+ax^{2}+bx^{2}+cx+d$ in the cubic extension $Q(\omega)$ and the result then follows from Theorem 1.
Case 2. $f(x,y)$ is reducible in $\mathbb{F}_{p}[x,y]$ where
@@
$I_{0}=[1,\frac{1}{4}p^{3/8}]$ and $K=[1,p^{\kappa}]$. We translate $(x,y)$ by $(stx_{0},sty_{0})$ with $x_{0},y_{0}\in I_{0}$ and $s,t\in K$ and estimate
@@
(15) $w(z_{1},z_{2}):=\\ \left|\left\{(x,y,x_{0},y_{0},s)\in I\times J\times I_{0}\times I_{0}\times K:z_{1}=\frac{x-\lambda_{1}y}{s(x_{0}-\lambda_{1}y_{0})},z_{2}=\frac{x-\omega_{2}y}{s(x_{0}-\omega_{2}y_{0})}\right\}\right|.\\$
@@
(16) $\frac{1}{|K|^{2}|I_{0}|^{2}}\sum\limits_{\begin{subarray}{c}z_{1}\in\mathbb{F}_{p}\\ z_{2}\in\mathbb{F}_{p^{2}}\end{subarray}}w(z_{1},z_{2})\left|\sum_{t\in K}\chi((t+z_{1})(t+z_{2})(t+z_{2}^{\prime}))\right|\\$
where $z_{2},z_{2}^{\prime}$ are the roots of a monic polynomial over $\mathbb{F}_{p}$. We shall use Holder’s inequality subsequently Weil’s Theorem to estimate the above. Using Holder’s inequality twice it is obtained
@@
The third term is estimated using Weil’s bound [9] and consequently by Lemma 11 as the following
$\left(\sum\limits_{\begin{subarray}{c}z_{1}\in\mathbb{F}_{p}\\ z_{2}\in\mathbb{F}_{p^{2}}\end{subarray}}\left|\sum_{t\in K}\chi((t+z_{1})(t+z_{2})(t+z_{2}^{\prime}))\right|^{2k}\right)^{\frac{1}{2k}}\ll\left(\sum\limits_{\begin{subarray}{c}z_{1}\in\mathbb{F}_{p}\\ a\in\mathbb{F}_{p}\\ b\in\mathbb{F}_{p}\end{subarray}}\left|\sum_{t\in K}\chi((t+z_{1})(t^{2}+a\,t+b))\right|^{2k}\right)^{\frac{1}{2k}}$
(19) $\ll r|K|^{\frac{1}{2}}p^{\frac{3}{2r}}+|K|p^{\frac{3}{4r}}\\$
@@
(20) $\sum w(z_{1},z_{2})^{2}\leq|I|^{2}|J|^{2}|K|^{2}p^{-\tau}$
for some $\tau>0$. Choosing $\kappa=3/2k$, we obtain an upper bound for the equation (14) of the form $|I||J|p^{-\tau/2k}$ from the above three estimates. For estimating second term we may remind the reader $|I|,|J|<p^{1/2}$ here. Let
@@
(21) $E(R,T,S)\leq|I|^{2}|J|^{2}|K|^{2}p^{-\tau}.$
First we estimate the following:
@@
$E(R,T)<\exp\left(c\frac{\log p}{\log\log p}\right).$
We shall prove the above lemma in due course. Let us start with writing $R$ as union of translates of a $T$ as bellow:
@@
$E(R,T)\leq\frac{|R|^{2}}{|T|^{2}}\max_{\zeta\in R}E(T+\zeta,T).$
Now, it is enough to show
$\max_{T+\zeta\,\subseteq\,R}E(T+\zeta,T)<\exp\left(c\frac{\log p}{\log\log p}\right)|T|^{2}.$
This is an immediate consequence of the Lemma below.
@@
$\max_{\rho\in\mathbb{F}_{p}^{*}\times\mathbb{F}_{p^{2}}^{*}\times\mathbb{F}_{p^{2}}^{*}}\left|\left\{(z_{1}+\zeta_{,}z_{2})\in T+\zeta\times T:\rho=z_{1}z_{2}\right\}\right|<\exp\left(c\frac{\log p}{\log\log p}\right)$
where, $T+\zeta\subseteq R$.
@@
Let us fix an $\rho=(\rho_{1},\rho_{2},\rho_{3})\in\mathbb{F}_{p}^{*}\times\mathbb{F}_{p^{2}}^{*}\times\mathbb{F}_{p^{2}}^{*}$ and consider a product representation of $\rho=(z_{1}+\zeta)\times z_{2}$ where $z_{1},z_{2}\in T$. Then $z_{1}=(x_{1}-\lambda_{1}y_{1},x_{1}-\omega_{2}y_{1},x_{1}-\omega_{3}y_{1})$, $z_{2}=(x_{2}-\lambda_{1}y_{2},x_{2}-\omega_{2}y_{2},x_{2}-\omega_{3}y_{2})$ and $\zeta=(x_{\zeta}-\lambda_{1}y_{\zeta},x_{\zeta}-\omega_{2}y_{\zeta},x_{\zeta}-\omega_{3}y_{\zeta})$ with $x_{1},x_{2},y_{1},y_{2}\in I_{0},x_{\zeta}\in I$ and $y_{\zeta}\in J$, we want to obtain an estimate of the number of solutions of the following:
(22) $\begin{cases}(x_{1}+x_{\zeta}-\lambda_{1}(y_{1}+y_{\zeta}))(x_{2}-\lambda_{1}y_{2})=\rho_{1}(\text{ in }F_{p}(\omega))\\ \\ (x_{1}+x_{\zeta}-\omega_{2}(y_{1}+y_{\zeta}))(x_{2}-\omega_{2}y_{2})=\rho_{2}(\text{ in }F_{p}(\omega))\\ \\ (x_{1}+x_{\zeta}-\omega_{3}(y_{1}+y_{\zeta}))(x_{2}-\omega_{3}y_{2})=\rho_{3}(\text{ in }F_{p}(\omega))\end{cases}$
where $\omega_{2},\omega_{3}$ are the root of the polynomial $x^{2}+\lambda_{2}x+\lambda_{3}.$ The estimate of number of solutions of (22) is again bounded by the number of solution of the following:
(23) $\begin{cases}(x_{1}-\lambda_{1}y_{1})(x_{2}-\lambda_{1}y_{2})=\rho_{1}(\text{ in }F_{p}(\omega))\\ \\ (x_{1}-\omega_{2}y_{1})(x_{2}-\omega_{2}y_{2})=\rho_{2}(\text{ in }F_{p}(\omega))\\ \\ (x_{1}-\omega_{3}y_{1})(x_{2}-\omega_{3}y_{2})=\rho_{3}(\text{ in }F_{p}(\omega))\end{cases}$
where, $x_{1}\in I,y_{1}\in J$ and $x_{2},y_{2}\in I_{0}$.
Let $\mathcal{S}$ be the solution set of the above equation. Let $(x_{1},x_{2},y_{1},y_{2}),(x_{1}^{\prime},x_{2}^{\prime},y_{1}^{\prime},y_{2}^{\prime})\in\mathcal{S}$. So, it satisfies the equation (22). So, all three of $\lambda_{1},\omega_{1},\omega_{3}$ satisfies the quadratic polynomial equation
@@
[…]
Also in the above congruences (24) the elements are in $\mathbb{F}_{p}$. So, this gives the same congruences in the field $\mathbb{F}_{p}$. Now by the definition of $I,J$ and $I_{0}$; their elements are bounded by $p^{1/2}$ and $\frac{p^{3/8}}{4}$ respectively. Therefore, $|x_{1}x_{2}-x_{1}^{\prime}x_{2}^{\prime}|,|y_{1}y_{2}-y_{1}^{\prime}y_{2}^{\prime}|$ are bounded by $p^{7/8}$. As both of them are elements of $\mathbb{Z}$ with an absolute value less than the modulus the equalities in (24) are, in fact, inequalities in $\mathbb{Z}$ with
@@
$y_{1}y_{2}=y_{1}^{\prime}y_{2}^{\prime}.$
Finally we shall again use the divisor estimate $\exp\left(c\frac{\log p}{\log\log p}\right)$ to get the required result. ∎
Now we shall continue with proving the equation (21). Let $\kappa=3/2k$ in the definition of $K$. Thus,
@@
(25) $E(R,T,S)\leq Q^{\frac{1}{k}}.\exp\left(c\frac{\log p}{\log\log p}\right).|R|^{2(1-\frac{1}{k})}$